OpenStudy (anonymous):

In 1980, median family income was about $18,000 and in 2000 it was about $43,000. Find the slope passing through points (1980,18000) and (2000,43000) m=? Family income in 2002?

OpenStudy (anonymous):

i have the answer but dont know if its right

OpenStudy (anonymous):

I have 43000-18000=25000 2000-1980=20 divide 25000/20= 1250

OpenStudy (anonymous):

i dont know what m=

OpenStudy (mathteacher1729):

m is exactly what you found when you divided 25,000 / 20 = 1,250 This is a VERY interesting question, and relevant, since you will one day have to pay your own bills... :p Play around with some of the choices. Compare the all-USA average with your state over time. Then look at the income vs. education and race... not everyone makes the same. Some of the differences might be surprising. http://www.google.com/publicdata/explore?ds=c8op9mhgodplq_&ctype=l&strail=false&bcs=d&nselm=h&met_y=median_income_current&scale_y=lin&ind_y=false&rdim=country&idim=country:US&ifdim=country&ind=false&icfg

OpenStudy (anonymous):

i got m=1250 and family income $1250